OpenStudy (anonymous):

4 sqrt 3 +2 sqrt 1/3

OpenStudy (anonymous):

Okay, so first we can break down the different terms. In the second term, we have\[\sqrt{1/3} = \sqrt{1/3} * \sqrt{3}/\sqrt{3} = \sqrt{3}/3,\] correct?

OpenStudy (anonymous):

yes and i thought u said u were doing it on a whiteboard. is that wat this is? (:

OpenStudy (anonymous):

so i have 4 sqrt 3 + 2 sqrt 3 over regular 3

OpenStudy (anonymous):

Nah, the link for the whiteboard site is elsewhere, I can give it to you if you'd like to go there though. :) Then, we can look at the whole: \[4 * \sqrt{3} + 2/3 * \sqrt{3}\] And we can factor out the sqrt(3): \[\sqrt{3}*(4+2/3)\] and you can solve that on your own. :)

OpenStudy (anonymous):

so in order to get common denominators the answer will turn out to be 14 sqrt 3/ regular 3 rightt...?

OpenStudy (anonymous):

Yep, you've got it. ;D

OpenStudy (anonymous):

thank u so much!!!!!